Shift report 20010814 08:00-16:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Tue Aug 14 2001 - 16:01:38 EDT

  • Next message: Flemming Videbaek: "RHIC near term plans"

    ------------------------------------------------------------
    Date: Tue Aug 14 16:01:37 2001, Shift: 08:00-16:00
    Supervisor: Andrei Makeev, with: 
    Log book#7 pages 87 - 89
    
    Shift Summary:
    ------------------------------------------------------------
    1. No beam all the time. RF and magnets problems
    at RHIC.
    2. The ZDC's high- and low-gain ADCs  gating has been
    corrected, since the cables were connected improperly.
    3.Si5_3 is not working; Steve and Hiro removed it from IR.
    
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Tue Aug 14 2001 - 16:02:30 EDT